OpenStudy (anonymous):

How do I solve: The Line with x-intercept -2 and y-intercept -1 in slope-intercept form?

OpenStudy (anonymous):

we got two points ( -2,0) (0,-1), which are the x-int and the y-int we know the slope form is m= (y^2-y^1)/(x^2-x^1) just plug in the numbers, you will find the slope

OpenStudy (whpalmer4):

If a line has an x-intercept of -2, that means the line goes through the point (-2,0). A line with a y-intercept of -1 goes through the point (0,-1). With the formula \[m = \frac{y_2-y_1}{x_2-x_1}\]we can find the slope \((m)\) of a line going through the two points \((x_1,y_1)\) and \((x_2,y_2)\) Then use the point-slope formula \[y-y_1 = m(x-x_1)\]to write an equation for the line, and solve for \(y\) to put it in slope-intercept form which is \[y = mx+b\]

OpenStudy (whpalmer4):

x^2 is typically understood to mean \(x^2\), not \(x_2\)